(D^2-2DD'+D^2)z=12xy
1
Expert's answer
2020-09-07T13:37:28-0400
"(D^2-2DD'+D'^2)z=12xy."

The auxiliary equation of the given equation is


"m^2-2m+1=0"

"(m-1)^2=0"

"m_{1,2}=1"

Then the complementary function of the given equation is


"u=\varphi_1(y+1,x)+x\varphi_2(y+1,x)=\varphi_1(y,x)+x\varphi_2(y,x)"

When "f(x, y)=V," where "V" is function of "x" and "y," then partial integral:


"p=\dfrac{1}{F(D,D')}V"

Then partial integral


"p=\dfrac{1}{D^2-2DD'+D'^2 }(12xy)="

"=\dfrac{1}{(D-D')^2 }(12xy)=\dfrac{12}{D^2 }(1-\dfrac{D'}{D})^{-2}(xy)="

"=\dfrac{12}{D^2 }(1+\dfrac{2D'}{D}+\dfrac{3D'^2}{D^2}+...)(xy)="

"=\dfrac{12}{D^2 }(xy+\dfrac{2}{D}(x)+0+...)="

"=\dfrac{12}{D^2 }(xy)+\dfrac{24}{D^3 }(x)=12y(\dfrac{x^3}{6 })+24(\dfrac{x^4}{24 })="


"=2x^3y+x^4"

Then the general solution of the given equation is "z=u+p," such that


"z=\varphi_1(y,x)+x\varphi_2(y,x)+x^4+2x^3y,"

where "\varphi_1, \varphi_2" are arbitrary functions.